OpenStudy (anonymous):

a^2 – 5a – 20 factor this trinomial i know how to do it but i cant figure it out , can sombody give me a strategy rather than guesses?

OpenStudy (amistre64):

complete the square

OpenStudy (anonymous):

Works every time....

OpenStudy (anonymous):

explain please?

OpenStudy (anonymous):

Formula is good as well...

OpenStudy (amistre64):

completing a square is a method where you create a perfect square binomial plus something extra

OpenStudy (anonymous):

Im still lost sorry?

OpenStudy (amistre64):

take half the "a" term and square it: (5/2)^2 add and subtract this from the original ... your essentially adding zero to it so that the value remains the same and only the form changes

OpenStudy (anonymous):

a^2 – 5a = 20 First off, write it like that...

OpenStudy (anonymous):

Ok, I will Amistre to it...:-)

OpenStudy (amistre64):

if you start out like estudier mentions, you add that value to each side to keep things balanced

OpenStudy (anonymous):

so a^2=20+5a?

OpenStudy (amistre64):

you might be better off just using the quadratic formula ... trying to expalin "completing the square" is abit difficult

OpenStudy (anonymous):

can you show me that then ?

OpenStudy (amistre64):

given a quadratic of the form:\[fx^2+mx+l\] we can use a formula to calculate the xs\[\frac{-m\pm\sqrt{m^2-4fl}}{2f}\]

OpenStudy (amistre64):

i woulda used the abcs, but it mighta confused you since your expression uses a as the variable

OpenStudy (anonymous):

can you plug in the numbers for me as an example?

OpenStudy (amistre64):

im sure youre capable of figureing out which number is the (f)irst, (m)iddle, and (l)ast to plug in ...

OpenStudy (anonymous):

Ok thanks

OpenStudy (amistre64):

if you want, i can chk your work to see if you did it correctly

OpenStudy (anonymous):

It has no rational zeros ...

OpenStudy (anonymous):

Correct solutions are irrational (that's why you can't easily guess at the roots).

OpenStudy (anonymous):

soooo...?

OpenStudy (anonymous):

The formula has given you the roots, right?